The total earnings of company A and company B was $9 500 000.The total earnings of company B and C was $5 450 000.Company A earnings was 6 times as much as company C earnings.How much did company B earn?
1
Expert's answer
2011-08-29T06:35:37-0400
Let's denote the total earnings of company A as X, of company B as Y, the company C as Z. Therefore we can construct the system of equations:
X + Y = 9 500 000
Y + Z = 5 450 000
X = 6Z

6Z + Y = 9,500,000
Z + Y = 5,450,000
5Z = 4, 050,000
Z = 810,000
X = 6Z = 4,860,000
Y = 5,450,000 - Z = 5,450,000 - 810,000 = 4,640,000
Answer: The total earnings of company A was $4,860,000, the total earnings of company B was $4,640,000, the total earnings of company C was $810,000.